OpenStudy (anonymous):

44 - 3 × (3 - 1)2 c: plz help.

OpenStudy (alfie):

Is c the velocity of light in a vacuum or you refer to: 44-3*(2)2x ?

OpenStudy (anonymous):

no.... i think its just pemdas....

OpenStudy (anonymous):

evaluate? use PEMDAS 44 - 3 × (3 - 1)2 given =44-3*(2)^2 inside Parentheses first =44-3*4 Exponents =44-12 Mutliplication =32 Subtraction

OpenStudy (katrinakaif):

PEDMAS 44 - 3 X (2) (2) 44 - 3 x 4 44 - 12

OpenStudy (anonymous):

thanx a bunch!